# Holography for the IKKT matrix model

### v1

@@ Abstract
Extrapolating this correspondence to the extremal case of $p=-1$ suggests a holographic duality between strings on the D$(-1)$ instanton background and super Yang-Mills theory reduced to zero dimensions. The latter is known as the IKKT matrix model, and has by itself been conjectured as a candidate for a nonperturbative formulation of type IIB superstring theory [3]. The dual D$(-1)$ instanton background is a half-supersymmetric solution of Euclidean IIB supergravity with flat spacetime metric and non-trivial dilaton/axion [4, 5, 6, 7]. Despite the compelling and highly non-trivial results that have been obtained in the IKKT model [8, 9, 10, 11], this holographic duality has remained largely unstudied, not at least due to various subtleties and degeneracies on both sides of the correspondence. However, recent work has started to probe holography in the polarized IKKT matrix model, a supersymmetric ${\rm SO}(3)\times{\rm SO}(7)$ invariant mass deformation of the original model [12, 13, 14].
@@
Let us start from the IKKT matrix model, obtained as the reduction of (Euclidean) ten-dimensional super Yang-Mills theory to zero dimensions, with the action given by
@@
$\delta_{\epsilon}\mathcal{O}^{ab}=\,$ $\tfrac{9}{5}\,\Gamma^{[a}\mathcal{O}^{b]}\,,$ (14)
@@
In this letter, we have constructed maximal supergravity in one dimension, finally completing the list of maximal supergravities across all spacetime dimensions. The model, by itself, stands out as a new and distinguished example of supersymmetric quantum mechanics, featuring 32 (complex) supercharges, and an intricate highly non-linear scalar potential which deserves further exploration. In particular, its quantization could be approached using techniques that bypass the technical difficulties of quantum field theory. It would also be interesting to explore whether the non-perturbative regime of the model can be studied using supersymmetric localization.
@@
We have also in this letter constructed the general 1/2-BPS solutions of the model, which all uplift to IIB solutions with flat spacetime metric, and preserve all the supersymmetries of the D$(-1)$ background. It will be interesting to explore in more detail potential relations to the known deformations of the IKKT model, in particular to the so-called polarized IKKT matrix model for which recent progress has been achieved in [12, 14, 13]. Another recently explored deformation is based on holography for the D($-1$)/D7 system [30] whose near-horizon background shares many similarities with our solutions [31].

### v2

@@ Abstract
Extrapolating this correspondence to the extremal case of $p=-1$ suggests a holographic duality between strings on the D$(-1)$ instanton background and super Yang-Mills theory reduced to zero dimensions. The latter is known as the IKKT matrix model, and has by itself been conjectured as a candidate for a nonperturbative formulation of type IIB superstring theory [3]. The dual D$(-1)$ instanton background is a half-supersymmetric solution of Euclidean IIB supergravity with flat spacetime metric and non-trivial dilaton/axion [4, 5, 6, 7]. Despite the compelling and highly non-trivial results that have been obtained in the IKKT model [8, 9, 10, 11], this holographic duality has remained largely unstudied, not at least due to various subtleties and degeneracies on both sides of the correspondence. However, recent work has started to probe holography in the polarized IKKT matrix model [12, 13, 14], a supersymmetric ${\rm SO}(3)\times{\rm SO}(7)$ invariant mass deformation of the original model [15].
@@
Let us start from the Euclidean IKKT matrix model, obtained as the reduction of ten-dimensional super Yang-Mills theory to zero dimensions, with the action given by
@@
$\delta_{\epsilon}\mathcal{O}^{ab}=\,$ $\tfrac{9}{5}\,\Gamma^{(a}\mathcal{O}^{b)}\,,$ (14)
@@
In this letter, we have constructed maximal supergravity in one dimension, finally completing the list of maximal supergravities across all spacetime dimensions. The model, by itself, stands out as a new and distinguished example of supersymmetric quantum mechanics, featuring 32 supercharges, and an intricate highly non-linear scalar potential which deserves further exploration. In particular, its quantization could be approached using techniques that bypass the technical difficulties of quantum field theory. It would also be interesting to explore whether the non-perturbative regime of the model can be studied using supersymmetric localization.
@@
We have also in this letter constructed the general 1/2-BPS solutions of the model, which all uplift to IIB solutions with flat spacetime metric, and preserve all the supersymmetries of the D$(-1)$ background. It will be interesting to explore in more detail potential relations to known deformations of the IKKT model, in particular to the so-called polarized IKKT matrix model [15] for which recent progress has been achieved in [12, 14, 13], as well as to the deformations that are dual to the spherical branes solutions of [31, 32]. Another recently explored deformation is based on holography for the D($-1$)/D7 system [33] whose near-horizon background shares many similarities with our solutions [34].